Bill drove from Ajax to Bixby at an average speed of 50mi/h. On the way back, he drove at 60mi/h. The total trip took 4 2/3 hours of driving time. Find the distance between the two cities.
Answer by checkley71(8403)

50T=60(14/3-T)
50T=20*14-60T
50T+60T=280
110T=280
T=280/110
T=2 60/110 OR 2 6/10 HOURS FOR THE 50 MPH TRIP.
4 2/3-2 6/10
14/3-26/10
(14*10-3*26)/30
(140-78)/30
62/30=2 1/15 HOURS FOR THE 60 MPH TRIP.
PROOF:
2 6/10+2 1/15
26/10+31/15
(26*15+31*10)/150
(390+310)/150
700/150= 4 2/3